For the reaction at 25°C, N2O4(g) = 2NO2(g) if delGf° for N2O4 and NO2 are 23.49 and 12.39 kcal respectively then Ko for the reaction is(in atm)...
Ans is 0.1132atm
 

Grade:12th pass

1 Answers

Arun
3 years ago
Let, 1 mol of N 
2
​ 
 O 
4
​ 
  present initially. 
 
                  N 
2
​ 
 O 
4
​ 
 ⇋2NO 
2
​ 
 
t=0: 1 0
t=t 
eq
​ 
 : 1−x 2x
 
So, the number of molecules at equilibrium =n 
2
​ 
 O 
4
​ 
 
​ 
 +n 
NO 
2
​ 
 
​ 
 =1+x−2x=1+x.